OpenStudy (anonymous):

Evaluate the definite integral (limits from 0 to 1) using substitution. x^13*(e^x^14) dx My question is which do you substitute?

OpenStudy (anonymous):

u = x^14

OpenStudy (anonymous):

ok so using u=x^14, i get \[1\div14\] \[\int\limits_{0}^{1}\] then x^13*e^(-u) du is the next step to evaluate using the limits?

OpenStudy (anonymous):

i cant seem to get all of the equation to be on the same line..anyways the 1/14 i isolated in front of the integral sign

OpenStudy (anonymous):

\[\int x^{13}e^{x^{14} }dx\]

OpenStudy (anonymous):

\[u=x^{14}\] \[du=14x^{13}dx\] \[\frac{du}{14}=x^{13}dx\] \[\int x^{13}e^{x^{14}}dx=\int e^{x^{14}}x^{13}dx=\int \frac{e^u}{14}du=\frac{1}{14}\int e^u du \]

OpenStudy (anonymous):

\[\int e^u=e^u+c\] \[\frac{1}{14}\int e^u du=\frac{1}{14}e^u+c=\frac{1}{14} e^{x^{14}}+c\]

OpenStudy (anonymous):

ok thank you i understand now :)